OpenStudy (anonymous):

Given that pi/2 is less than or equal to theta, 3pi/2 is greater than or equal to theta, and tan of theta is 3, what is the value of cos2theta?

OpenStudy (anonymous):

\[\pi/2\le \emptyset \le3\pi/2 and \tan \emptyset=3, find the value of \cos2\]

OpenStudy (anonymous):

Can anyone tell me how to find the sine and cosine of tan theta=3? I know the rest

OpenStudy (anonymous):

instead of theta i am using x given tanx=3 find cos(2x) now we know cos(2x)=(1-(tanx)^2)/(1+(tanx)^2)=(1-9)(1+9)=-8/10=-4/5

OpenStudy (anonymous):

if tanx =3 then sec^2x=1+tan^2x=1+9=10 hence cosx= 1/secx =1/sqrt(10) (use proper sign after looking at the range of x) also sin^2x=1-cos^2x=1-1/10 hence sinx =3/sqrt(10) (use proper sign after looking at the range of x)